OpenStudy (anonymous):

Calculate the mass defect (in amu) for 23Na (which has the atomic mass 22.9898 amu). Mass of 1H atom = 1.0078 amu Mass of neutron = 1.0087 amu Mass of proton = 1.0073 amu Mass of electron = 5.4858×10-4 amu amu You do not necessarily need to use all of the masses given above. Calculate the binding energy per nucleon (in MeV) for the 23Na nucleus. MeV

OpenStudy (aaronq):

The mass defect is the apparent loss of mass when the individual pieces (e.g. protons, neutrons) to form atoms. Mass Defect = mass of all protons, neutrons, electrons - atomic mass Sodium has 11 protons, and therefore 23Na has 12 neutrons. There are 11 electrons as well. Mass defect=1.0073*11 + 1.0087*12+11*(\(5.4858×10^{-4})\)- 22.9898 =0.20093438 amu Convert them to kilograms. Use in \(\sf E=mc^2\) to find the energy (which will be in joules). You'll also need to convert to MeV, \(\sf 1~MeV=1.6*10^{-13}~J\)